OpenStudy (danaplague):

(01.05) What is the equation in point-slope form of the line passing through (0, 5) and (−2, 11)? y − 5 = −3(x + 2) y − 5 = 3(x + 2) y − 11 = −3(x − 2) y − 11 = −3(x + 2)

OpenStudy (jiteshmeghwal9):

Slope = (11-5)/(-2-0)=-3

OpenStudy (danaplague):

thankssss

OpenStudy (jiteshmeghwal9):

Equation format (Y-y_1)=m(x-x_1) Where x_1=0 Y_1=5 m=-3

OpenStudy (3mar):

In other words,\[\frac{ y-y_{1} }{ x-x_{1} }=\frac{ y_2-y_{1} }{ x_2-x_{1} }\]. Thanks, jiteshmeghwal9. You deserve a medal! \[\frac{ y-5 }{ x-0 }=\frac{ 11-5 }{ -2-0 }\] \[\frac{ y-5 }{ x }=-3\] \[y-5=-3x\] \[y=-3x+5\]

OpenStudy (jiteshmeghwal9):

Thanx for appreciation @3mar

OpenStudy (3mar):

Not at all, brother!